# A Genealogy of Multi-Sensor Foundation Models in Remote Sensing

### v1

@@ Acknowledgements.
- Wang et al. (2022) Yi Wang, Conrad M. Albrecht, Nassim Ait Ali Braham, Lichao Mou, and Xiao Xiang Zhu. 2022. Self-Supervised Learning in Remote Sensing: A review. IEEE Geoscience and Remote Sensing Magazine 10, 4 (Dec. 2022), 213–247. doi:10.1109/MGRS.2022.3198244 Conference Name: IEEE Geoscience and Remote Sensing Magazine.
@@
- Wang et al. (2023a) Yi Wang, Hugo Hernández Hernández, Conrad M. Albrecht, and Xiao Xiang Zhu. 2023a. Feature Guided Masked Autoencoder for Self-supervised Learning in Remote Sensing. doi:10.48550/arXiv.2310.18653 arXiv:2310.18653 [cs].
@@
- Werninghaus and Buckreuss (2010) Rolf Werninghaus and Stefan Buckreuss. 2010. The TerraSAR-X Mission and System Design. IEEE Transactions on Geoscience and Remote Sensing 48, 2 (Feb. 2010), 606–614. doi:10.1109/TGRS.2009.2031062
@@
- Wu et al. (2024) Zhiyu Wu, Xiaokang Chen, Zizheng Pan, Xingchao Liu, Wen Liu, Damai Dai, Huazuo Gao, Yiyang Ma, Chengyue Wu, Bingxuan Wang, Zhenda Xie, Yu Wu, Kai Hu, Jiawei Wang, Yaofeng Sun, Yukun Li, Yishi Piao, Kang Guan, Aixin Liu, Xin Xie, Yuxiang You, Kai Dong, Xingkai Yu, Haowei Zhang, Liang Zhao, Yisong Wang, and Chong Ruan. 2024. DeepSeek-VL2: Mixture-of-Experts Vision-Language Models for Advanced Multimodal Understanding. doi:10.48550/arXiv.2412.10302 arXiv:2412.10302 [cs].
@@
- Wulder et al. (2022) Michael A. Wulder, David P. Roy, Volker C. Radeloff, Thomas R. Loveland, Martha C. Anderson, David M. Johnson, Sean Healey, Zhe Zhu, Theodore A. Scambos, Nima Pahlevan, Matthew Hansen, Noel Gorelick, Christopher J. Crawford, Jeffrey G. Masek, Txomin Hermosilla, Joanne C. White, Alan S. Belward, Crystal Schaaf, Curtis E. Woodcock, Justin L. Huntington, Leo Lymburner, Patrick Hostert, Feng Gao, Alexei Lyapustin, Jean-Francois Pekel, Peter Strobl, and Bruce D. Cook. 2022. Fifty years of Landsat science and impacts. Remote Sensing of Environment 280 (Oct. 2022), 113195. doi:10.1016/j.rse.2022.113195
@@
- Xiong et al. (2024) Zhitong Xiong, Yi Wang, Fahong Zhang, Adam J. Stewart, Joëlle Hanna, Damian Borth, Ioannis Papoutsis, Bertrand Le Saux, Gustau Camps-Valls, and Xiao Xiang Zhu. 2024. Neural Plasticity-Inspired Multimodal Foundation Model for Earth Observation. doi:10.48550/arXiv.2403.15356 arXiv:2403.15356 [cs].
@@
- Zhang et al. (2022) Qiming Zhang, Yufei Xu, Jing Zhang, and Dacheng Tao. 2022. VSA: Learning Varied-Size Window Attention in Vision Transformers. In Computer Vision – ECCV 2022, Shai Avidan, Gabriel Brostow, Moustapha Cissé, Giovanni Maria Farinella, and Tal Hassner (Eds.). Springer Nature Switzerland, Cham, 466–483. doi:10.1007/978-3-031-19806-9_27
- Zhang et al. (2024) Weijia Zhang, Jindong Han, Zhao Xu, Hang Ni, Hao Liu, and Hui Xiong. 2024. Urban Foundation Models: A Survey. In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24). Association for Computing Machinery, New York, NY, USA, 6633–6643. doi:10.1145/3637528.3671453
- Zhao et al. (2016) Lijun Zhao, Ping Tang, and Lianzhi Huo. 2016. Feature significance-based multibag-of-visual-words model for remote sensing image scene classification. Journal of Applied Remote Sensing 10, 3 (July 2016), 035004. doi:10.1117/1.JRS.10.035004 Publisher: SPIE.
- Zhou et al. (2024) Yue Zhou, Litong Feng, Yiping Ke, Xue Jiang, Junchi Yan, Xue Yang, and Wayne Zhang. 2024. Towards Vision-Language Geo-Foundation Model: A Survey. doi:10.48550/arXiv.2406.09385 arXiv:2406.09385 [cs].
- Zhu et al. (2019) Xiao Xiang Zhu, Jingliang Hu, Chunping Qiu, Yilei Shi, Jian Kang, Lichao Mou, Hossein Bagheri, Matthias Häberle, Yuansheng Hua, Rong Huang, Lloyd Hughes, Hao Li, Yao Sun, Guichen Zhang, Shiyao Han, Michael Schmitt, and Yuanyuan Wang. 2019. So2Sat LCZ42: A Benchmark Dataset for Global Local Climate Zones Classification. doi:10.48550/arXiv.1912.12171 arXiv:1912.12171 [cs].

### v2

@@ Acknowledgements.
- Wang et al. (2022b) Yi Wang, Conrad M. Albrecht, Nassim Ait Ali Braham, Lichao Mou, and Xiao Xiang Zhu. 2022b. Self-Supervised Learning in Remote Sensing: A review. IEEE Geoscience and Remote Sensing Magazine 10, 4 (Dec. 2022), 213–247. https://doi.org/10.1109/MGRS.2022.3198244 Conference Name: IEEE Geoscience and Remote Sensing Magazine.
- Wang et al. (2022a) Yi Wang, Conrad M Albrecht, and Xiao Xiang Zhu. 2022a. Self-Supervised Vision Transformers for Joint SAR-Optical Representation Learning. In IGARSS 2022 - 2022 IEEE International Geoscience and Remote Sensing Symposium. 139–142. https://doi.org/10.1109/IGARSS46834.2022.9883983 ISSN: 2153-7003.
- Wang et al. (2023a) Yi Wang, Nassim Ait Ali Braham, Zhitong Xiong, Chenying Liu, Conrad M. Albrecht, and Xiao Xiang Zhu. 2023a. SSL4EO-S12: A Large-Scale Multi-Modal, Multi-Temporal Dataset for Self-Supervised Learning in Earth Observation. https://doi.org/10.48550/arXiv.2211.07044 arXiv:2211.07044.
@@
- Wang et al. (2023b) Yi Wang, Hugo Hernández Hernández, Conrad M. Albrecht, and Xiao Xiang Zhu. 2023b. Feature Guided Masked Autoencoder for Self-supervised Learning in Remote Sensing. https://doi.org/10.48550/arXiv.2310.18653 arXiv:2310.18653 [cs].
@@
- Werninghaus and Buckreuss (2010) Rolf Werninghaus and Stefan Buckreuss. 2010. The TerraSAR-X Mission and System Design. IEEE Transactions on Geoscience and Remote Sensing 48, 2 (Feb. 2010), 606–614. https://doi.org/10.1109/TGRS.2009.2031062
@@
- Wu et al. (2024) Zhiyu Wu, Xiaokang Chen, Zizheng Pan, Xingchao Liu, Wen Liu, Damai Dai, Huazuo Gao, Yiyang Ma, Chengyue Wu, Bingxuan Wang, Zhenda Xie, Yu Wu, Kai Hu, Jiawei Wang, Yaofeng Sun, Yukun Li, Yishi Piao, Kang Guan, Aixin Liu, Xin Xie, Yuxiang You, Kai Dong, Xingkai Yu, Haowei Zhang, Liang Zhao, Yisong Wang, and Chong Ruan. 2024. DeepSeek-VL2: Mixture-of-Experts Vision-Language Models for Advanced Multimodal Understanding. https://doi.org/10.48550/arXiv.2412.10302 arXiv:2412.10302 [cs].
@@
- Wulder et al. (2022) Michael A. Wulder, David P. Roy, Volker C. Radeloff, Thomas R. Loveland, Martha C. Anderson, David M. Johnson, Sean Healey, Zhe Zhu, Theodore A. Scambos, Nima Pahlevan, Matthew Hansen, Noel Gorelick, Christopher J. Crawford, Jeffrey G. Masek, Txomin Hermosilla, Joanne C. White, Alan S. Belward, Crystal Schaaf, Curtis E. Woodcock, Justin L. Huntington, Leo Lymburner, Patrick Hostert, Feng Gao, Alexei Lyapustin, Jean-Francois Pekel, Peter Strobl, and Bruce D. Cook. 2022. Fifty years of Landsat science and impacts. Remote Sensing of Environment 280 (Oct. 2022), 113195. https://doi.org/10.1016/j.rse.2022.113195
- Xie et al. (2022) Zhenda Xie, Zheng Zhang, Yue Cao, Yutong Lin, Jianmin Bao, Zhuliang Yao, Qi Dai, and Han Hu. 2022. SimMIM: A Simple Framework for Masked Image Modeling. 9653–9663. https://openaccess.thecvf.com/content/CVPR2022/html/Xie_SimMIM_A_Simple_Framework_for_Masked_Image_Modeling_CVPR_2022_paper.html
- Xie et al. (2023) Zhenda Xie, Zheng Zhang, Yue Cao, Yutong Lin, Yixuan Wei, Qi Dai, and Han Hu. 2023. On Data Scaling in Masked Image Modeling. 10365–10374. https://openaccess.thecvf.com/content/CVPR2023/html/Xie_On_Data_Scaling_in_Masked_Image_Modeling_CVPR_2023_paper.html
@@
- Xiong et al. (2024) Zhitong Xiong, Yi Wang, Fahong Zhang, Adam J. Stewart, Joëlle Hanna, Damian Borth, Ioannis Papoutsis, Bertrand Le Saux, Gustau Camps-Valls, and Xiao Xiang Zhu. 2024. Neural Plasticity-Inspired Multimodal Foundation Model for Earth Observation. https://doi.org/10.48550/arXiv.2403.15356 arXiv:2403.15356 [cs].
@@
- Yıldız et al. (2025) Çağatay Yıldız, Nishaanth Kanna Ravichandran, Nitin Sharma, Matthias Bethge, and Beyza Ermis. 2025. Investigating Continual Pretraining in Large Language Models: Insights and Implications. https://doi.org/10.48550/arXiv.2402.17400 arXiv:2402.17400 [cs].
- Zbontar et al. (2021) Jure Zbontar, Li Jing, Ishan Misra, Yann LeCun, and Stephane Deny. 2021. Barlow Twins: Self-Supervised Learning via Redundancy Reduction. In Proceedings of the 38th International Conference on Machine Learning. PMLR, 12310–12320. https://proceedings.mlr.press/v139/zbontar21a.html ISSN: 2640-3498.
- Zhang et al. (2022) Qiming Zhang, Yufei Xu, Jing Zhang, and Dacheng Tao. 2022. VSA: Learning Varied-Size Window Attention in Vision Transformers. In Computer Vision – ECCV 2022, Shai Avidan, Gabriel Brostow, Moustapha Cissé, Giovanni Maria Farinella, and Tal Hassner (Eds.). Springer Nature Switzerland, Cham, 466–483. https://doi.org/10.1007/978-3-031-19806-9_27
- Zhang et al. (2024) Weijia Zhang, Jindong Han, Zhao Xu, Hang Ni, Hao Liu, and Hui Xiong. 2024. Urban Foundation Models: A Survey. In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24). Association for Computing Machinery, New York, NY, USA, 6633–6643. https://doi.org/10.1145/3637528.3671453
- Zhao et al. (2016) Lijun Zhao, Ping Tang, and Lianzhi Huo. 2016. Feature significance-based multibag-of-visual-words model for remote sensing image scene classification. Journal of Applied Remote Sensing 10, 3 (July 2016), 035004. https://doi.org/10.1117/1.JRS.10.035004 Publisher: SPIE.
- Zhou et al. (2022) Jinghao Zhou, Chen Wei, Huiyu Wang, Wei Shen, Cihang Xie, Alan Yuille, and Tao Kong. 2022. iBOT: Image BERT Pre-Training with Online Tokenizer. https://doi.org/10.48550/arXiv.2111.07832 arXiv:2111.07832 [cs].
- Zhou et al. (2024) Yue Zhou, Litong Feng, Yiping Ke, Xue Jiang, Junchi Yan, Xue Yang, and Wayne Zhang. 2024. Towards Vision-Language Geo-Foundation Model: A Survey. https://doi.org/10.48550/arXiv.2406.09385 arXiv:2406.09385 [cs].
- Zhu et al. (2019) Xiao Xiang Zhu, Jingliang Hu, Chunping Qiu, Yilei Shi, Jian Kang, Lichao Mou, Hossein Bagheri, Matthias Häberle, Yuansheng Hua, Rong Huang, Lloyd Hughes, Hao Li, Yao Sun, Guichen Zhang, Shiyao Han, Michael Schmitt, and Yuanyuan Wang. 2019. So2Sat LCZ42: A Benchmark Dataset for Global Local Climate Zones Classification. https://doi.org/10.48550/arXiv.1912.12171 arXiv:1912.12171 [cs].